OpenStudy (anonymous):

What is the slope of the line that goes through (6, 5) and (3, 1)? The slope of the line that goes through (−3, −5) and (−3, −6) is ______.

OpenStudy (anonymous):

A. 3/4 B. -3/4 C. 4/3 D.-4/3

OpenStudy (anonymous):

A.positive B.negative C.zero D.undefined

OpenStudy (anonymous):

slope of line that goes through (6,5) and (3,1) slope = y2 - y1/x2 - x1 slope = 1 - 5/3 - 6 slope = - 4/-3 slope = 4/3 slope of a line that goes through (-3,-5) and (-3,-6) slope = y2 - y1/x2 - x1 slope = -6 - (-5)/-3 - (-3) slope = -6 + 5 / -3 + 3 slope = -1/0 = undefined

OpenStudy (anonymous):

Thank you so very much :))